Factorise the following expressions.
(i) 7x – 42
(ii) 6p – 12q
(iii) 7a2 + 14a
(iv) -16z + 20z3
(v) 20l2m + 30alm
(vi) 5x2y – 15xy2
(vii) 10a2 – 15b2 + 20c2
(viii) -4a2 + 4ab – 4ca
(ix) x2yz + xy2z + xyz2
(x) ax2y + bxy2 + cxyz

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
(i) 7x – 42 = 7(x – 6) (ii) 6p – 12q = 6(p – 2q) (iii) 7a2 + 14a = 7a(a + 2) (iv) -16z + 20z3 = 4z(-4 + 5z2) (v) 20l2m + 30alm = 10lm(2l + 3a) (vi) 5x2y – 15xy2 = 5xy(x – 3y) (vii) 10a2 – 15b2 + 20c2 = 5(2a2 – 3b2 + 4c2) (viii) -4a2 + 4ab – 4ca = 4a(-a + b – c) (ix) x2yz + xy2z + xyz2 = xyz(x + y + z) (x) ax2y + bxy2 + cxyz = xy(ax + by + cz)